OpenStudy (anonymous):

solve for x : 1000=500 e^2x a. x=ln(2)-2 b. x=ln(1000)/2*ln(500) c. x=ln(2)/2 d. x=1/e

OpenStudy (jdoe0001):

$$\bf 1000=500 e^2x\\ \frac{1000}{500}= e^{2x} \implies 2 = e^{2x}\\ \text{using log cancellation rule of }\\ log_aa^x = x\\ \text{keep in mind that }ln \implies log_e\\ log_e(2) = log_e(e^{2x})\\ log_e(2) = 2x $$

OpenStudy (jdoe0001):

or 2x = ln(2)

OpenStudy (anonymous):

so the answer would be C?